OpenStudy (wkep1234):

Use the Law of Sines to find the missing angle of the triangle. Find the measure of

OpenStudy (aaronandyson):

Just substitute the values in the law of sines.

OpenStudy (aaronandyson):

sin(A)/a = sin(C)/c plug in the given info to get sin(31)/44 = sin(C)/83 now solve for angle C sin(31)/44 = sin(C)/83 83*sin(31)/44 = sin(C) 0.97154909585 = sin(C) ... use a calculator here sin(C) = 0.97154909585 C = arcsin(0.97154909585) C = 76.2999941 ... use a calculator here C = 76.3 Now use this to find angle B A+B+C = 180 31 + B + 76.3 = 180 B + 107.3 = 180 B = 180 - 107.3 B = 72.7 So angle B is 72.7 degrees